Grade 12Integral Calculus

∫ecos^2x.sin2x dx

Lovey
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int e^{cos^{2}x}sin2x.dx
cos^{2}x = t
2cosx(-sinx)dx = dt
-sin2xdx = dt
I = \int e^{t}dt
I = e^{t}+c
I = e^{cos^{2}x}+c